# The COSINUS Collaboration

### v1

@@ IV Effect of energy resolution and quenching factor
$\sigma_{\rm DAMA}(QF_{T}E_{\rm nr})=(a\ {\rm{keV}_{ee}})\sqrt{QF_{T}E_{\rm nr}/{\rm{keV}_{ee}}}+bQF_{T}E_{\rm nr},$ (9)
@@
where the nominal values for the parameters are $a=0.448$, $b=0.0091$ [24]. In figure 4 we show how the required exposure for COSINUS depends on these parameters, using 3 (left) or 5 (right) bins in the unfolding procedure, with $N_{\rm iter}=5$ and $E_{\rm min}=1$ keV. These results were obtained by repeating the analysis procedure described in section III for varying values of the resolution function parameters $a$ and $b$. We observe that the required exposure is rather insensitive to the assumed DAMA/LIBRA resolution, so that the conclusions of the above analysis hold even if there is some uncertainty on the true energy resolution.
@@
Figure 4: The required exposure for 90% CL rejection of the DAMA/LIBRA dark matter signal, for $E_{\rm min}=1$ keV, $N_{\rm bin}=3,5$ (left, right) and $N_{\rm iter}=5$, as a function of the resolution function parameters $a,b$ in the parametrization (9). The nominal values for the parameters are shown by the black dot.
@@
To account for a possibly energy dependent quenching factor for ${\rm Na}$ in DAMA/LIBRA we use a parametrization motivated by the Lindhard model [25]
$QF_{\rm Na}(E_{\rm nr})=\beta\frac{\alpha g(E_{\rm nr})}{1+\alpha g(E_{\rm nr})},\quad g=3E_{\rm nr}^{0.15}+0.7E_{\rm nr}^{0.6}+E_{\rm nr}.$ (10)
@@
A fit of this form to a sample of measured values for the quenching factor [26, 27, 28] is shown in figure 5. We only consider the energy dependence of the ${\rm Na}$ $QF$ here, since the scenario where most of the scattering is off Iodine is in any case more favorable for COSINUS, as is evident from table 1.
@@
Figure 5: A fit to the quenching factor data [26, 27, 28] with the parametric form (10), with $\alpha=0.238$, $\beta=0.20$.
@@
We then vary the parameters $\alpha,\beta$ in (10) and repeat the analysis procedure described in section III with $N_{\rm iter}=5$, $E_{\rm min}=1$ keV. The resulting required exposure for a 90% CL exclusion of the DAMA/LIBRA DM signal is shown in figure 6. We again notice that the result is not very sensitive to changes in the quenching factor for realistic values of the Lindhard model parameters. Therefore our results can be deemed reliable regardless of the uncertainty about the DAMA/LIBRA quenching factors.
@@
Figure 6: The required exposure for 90% CL rejection of the DAMA/LIBRA dark matter signal, for $E_{\rm min}=1$ keV, $N_{\rm bin}=3,5$ (left, right) and $N_{\rm iter}=5$, as a function of the quenching factor parameters $\alpha,\beta$ using the parametrization (10). The best fit values for the parameters are shown by the black dot. The DAMA/LIBRA nominal values correspond to $\alpha=\infty$, $\beta=0.3$.
@@
This insensitivity to the variation in the $QF$ is largely due to the assumed 1 keV energy threshold and zero background above the threshold. Therefore we also comment on a hypothetical worst case scenario, where COSINUS would be background dominated in the low energy region, so that the effective event selection threshold would be as high as 4 keV. If simultaneously the DAMA/LIBRA quenching factor would be abnormally high, this could lead to a large portion of the DAMA/LIBRA signal being hidden below the COSINUS threshold. In an extreme scenario with $QF_{\rm Na}=1$, $E_{\rm min}=4\ {\rm keV}$ we obtain the required exposure as 240 kg day, 700 kg day and 2300 kg day for $1\sigma$, $2\sigma$ and $3\sigma$ confidence level exclusion, respectively. This, although unrealistic, worse case scenario serves to highlight the importance of reaching a sufficiently low energy threshold in COSINUS, to be able to see the DAMA/LIBRA signal even in the case of an abnormally high quenching factor hypothesis.
V Conclusions
@@
In the context of a given dark matter model, the results of any two experiments can be compared by fitting the model to both data sets, and evaluating the goodness of fit to the combination of data. However, for a model-independent analysis such forward modeling is not possible. Therefore we have performed an unfolding of the dark matter event rate implied by the DAMA/LIBRA annual modulation data in order to estimate the sensitivity of the COSINUS experiment to such event rate. We find that in an optimistic zero-background scenario and a 1 keV nuclear recoil detection threshold, COSINUS can reach a $1\sigma$ exclusion of the DAMA/LIBRA DM signal with 57 kg day exposure. For $2\sigma$ exclusion the required exposure is 130 kg day and for $3\sigma$ exclusion 250 kg day. Assuming gaussian errors for the unfolded distributions we estimate the required exposure for a $5\sigma$ exclusion as 750 kg day. We have investigated the sensitivity of these results to assumptions made about the DAMA/LIBRA quenching factor and energy resolution, and found the results to be stable under reasonable variations in these parameters.
Acknowledgements.

### v2

@@ 4 Effect of energy resolution and quenching factor
$\sigma_{\rm DAMA}(Q\!F_{T}E_{\rm nr})=(a\ {\rm{keV}_{ee}})\sqrt{Q\!F_{T}E_{\rm nr}/{\rm{keV}_{ee}}}+bQ\!F_{T}E_{\rm nr},$ (4.1)
@@
where the nominal values for the parameters are $a=0.448$, $b=0.0091$ [29]. In figure 5, we show how the required exposure for COSINUS depends on these parameters, using 3 (left) or 5 (right) bins in the unfolding procedure, with $N_{\rm iter}=5$ and $E_{\rm min}=1$ keV. These results were obtained by repeating the analysis procedure described in section 3 for varying values of the resolution function parameters $a$ and $b$. We observe that the required exposure is rather insensitive to the assumed DAMA/LIBRA resolution, varying by $\sim\pm 10\%$ within the scanned parameter range, so that the conclusions of the above analysis hold even if there is some uncertainty on the true energy resolution.
@@
Figure 5: The required exposure for 90% CL rejection of the DAMA/LIBRA dark matter signal, for $E_{\rm min}=1$ keV, $N_{\rm bin}=3,5$ (left, right) and $N_{\rm iter}=5$, as a function of the resolution function parameters $a,b$ in the parametrization (4.1). The nominal values for the parameters are shown by the black dot.
@@
To account for a possibly energy-dependent quenching factor for $\rm Na$} in DAMA/LIBRA, we use a parametrization motivated by the Lindhard model [31]:
$Q\!F_{\rm Na}(E_{\rm nr})=\beta\frac{\alpha g(E_{\rm nr})}{1+\alpha g(E_{\rm nr})},\quad g=3E_{\rm nr}^{0.15}+0.7E_{\rm nr}^{0.6}+E_{\rm nr}.$ (4.2)
@@
Here, the parameters $\alpha$ and $\beta$ control the shape and normalization of the QFQ\!F, respectively. A fit of this form to a sample of measured values for the quenching factor [32, 33, 34, 35, 16, 36] is shown in figure 6. We observe that this parametrization is not a very good fit to the data, with the best-fit $\chi^{2}/{\rm dof.}=5.6$. However, the purpose of this exercise is simply to characterize the dependence of our results on the possibility that the DAMA/LIBRA QFQ\!Fs might differ from the constant values used in the analysis, and could depend on the recoil energy in some way. Therefore the particular form of the energy dependence is not our main interest here. We only consider the energy dependence of the $\rm Na$} QFQ\!F here, since the scenario in which most of the scattering is off Iodine is, in any case, more favorable for COSINUS, as is evident from table 1.
@@
Figure 6: A fit with $\chi^{2}/{\rm dof.}=5.6$ to the quenching factor data [32, 33, 34, 35, 16, 36] with the parametric form (4.2), with $\alpha=0.213$, $\beta=0.206$.
@@
We then vary the parameters $\alpha,\beta$ in eq. (4.2) and repeat the analysis procedure described in section 3 with $N_{\rm iter}=5$, $E_{\rm min}=1$ keV. The resulting required exposure for a 90% CL exclusion of the DAMA/LIBRA DM signal is shown in figure 7. We again notice that the result is not very sensitive to changes in the quenching factor for realistic values of the Lindhard model parameters. Therefore our results can be deemed reliable regardless of the uncertainty about the DAMA/LIBRA quenching factors.
@@
Figure 7: The required exposure for 90% CL rejection of the DAMA/LIBRA dark matter signal, for $E_{\rm min}=1$ keV, $N_{\rm bin}=3,5$ (left, right) and $N_{\rm iter}=5$, as a function of the quenching factor parameters $\alpha,\beta$ using the parametrization (4.2). The best fit values for the parameters are shown by the black dot. The DAMA/LIBRA nominal values correspond to $\alpha=\infty$, $\beta=0.3$.
@@
This insensitivity to the variation in the QFQ\!F is largely due to the assumed 1 keV energy threshold and zero background above the threshold. Therefore, we also comment on a hypothetical worst case scenario, where COSINUS would be background dominated in the low energy region, so that the effective event selection threshold would be as high as 4 keV. If, simultaneously, the DAMA/LIBRA quenching factor were abnormally high, this could lead to a large portion of the DAMA/LIBRA signal being hidden below the COSINUS threshold. In an extreme scenario with $Q\!F_{\rm Na}=1$, $E_{\rm min}=4\ {\rm keV}$, we obtain the required exposure to be 240 kg day, 700 kg day and 2300 kg day for $1\sigma$, $2\sigma$ and $3\sigma$ confidence level exclusions, respectively. Although unrealistic, this worse case scenario serves to highlight the importance of reaching a sufficiently low energy threshold in COSINUS, to be able to see the DAMA/LIBRA signal even in the case of an abnormally high quenching factor hypothesis.
5 Conclusions
@@
In the context of a given dark matter model, the results of any two experiments can be compared by fitting the model to both data sets, and evaluating the goodness of fit to the combination of data. However, for a model-independent analysis, such forward modeling is not possible. Therefore, we have performed an unfolding of the dark matter event rate implied by the DAMA/LIBRA annual modulation data in order to estimate the sensitivity of the COSINUS experiment to such event rate. We find that, in an optimistic zero-background scenario and a 1 keV nuclear recoil detection threshold, COSINUS can reach a $1\sigma$ exclusion of the DAMA/LIBRA DM signal with 57 kg day exposure. For $2\sigma$ exclusion the required exposure is 130 kg day, and for $3\sigma$ exclusion, 250 kg day. Assuming gaussian errors for the unfolded distributions, we estimate the required exposure for a $5\sigma$ exclusion to be 750 kg day. A simple background model indicates that the low energy region below $\sim 3.5$ keV could be contaminated by electron recoil events leaking into the nuclear recoil selection window. If the event selection is restricted above 4 keV, we find the required exposure as 77 kg day ($1\sigma$), 170 kg day ($2\sigma$) and 320 kg day ($3\sigma$). We have investigated the sensitivity of these results to assumptions made about the DAMA/LIBRA quenching factor and energy resolution, and found them to be stable under reasonable variations in these parameters.
Acknowledgements